Rita rows 12 km upstream and 12 km downstream in 3 hours. The speed of her boat in still water is 9 km/hr. Find the speed of the stream
Answer by stanbon(75887)

Rita rows 12 km upstream and 12 km downstream in 3 hours. The speed of her boat in still water is 9 km/hr. Find the speed of the stream
----------------------
Let speed of the stream be "s"
-----------------------
Upstream DATA;
distance = 12 km ; Rate = 9-s km/h; Time = d/r = 12/(9-s) hrs.
-----------------------
Downstream DATA:
distance = 12 km ; Rate = 9+s; Time = 12/(9+s) hrs.
-----------------------
EQUATION:
time up + time down = 3 hrs
12/(9-s) + 12/(9+s) = 3
1/(9-s) + 1/(9+s) = 1/4
4(9+s) + 4(9-s) = 9^2-s^2
72 = 81-s^2
x^2 = 9
s = 3
===========
Cheers,
Stan H.